# A Tale of Tails: Model Collapse as a Change of Scaling Laws

### v1

@@ Appendix A Prior Work
demonstrating “emergence” of good BLEU performance with scale. This type of “emergence” has been massively confirmed in (Wei et al., 2022), where a working definition of “emerging” is “not present in smaller models, but appears in larger models”. In this sense, (Wei et al., 2022) demonstrate empirically a large number of “skills” appearing with scale, like Multi-Task NLU, Modular arithmetic, word unscrambling and transliteration.
@@
Strengthening the tie between scaling laws and emergent skill, albeit in the opposite direction, (Michaud et al., 2023) posit that skills that emerge in ”quanta” imply a scaling law of the loss. Related, (Chen et al., 2023) assume a hierarchy of skills to derive data curation mechanisms to precipitate the emergence of skills, though they do not allude to scaling laws directly.
@@ Appendix B Complimentary Figures for Sections 2, 3 and 4
Both Figures 3 (middle) and 12 illustrate the setting where each model consumes as much training data as its predecessor ($T_{0}=T$). We now relax the assumption that each successive model has strictly the same amount of training data as its predecessor. We assume that the generation 0 model is trained on $T_{0}$ (here, $T_{0}=100,000$) amount of original data to generate AI data for generation 1. All future generations, starting from generation 2, are trained on data generated by the most powerful model from the previous generation ($T=1,000,000$ data in this case). Figure 14 (for Hutter LLM) and 14 (for Hutter++ on paired bigram data) show the resulting scaling behavior. We take this setting even further by adding a top-p tail cutting mechanism and a temperature scaling mechanism for each synthetic data generation. Figure 6 cuts at $p=0.95$ and Figure 16 at temperature $0.9$.
@@
Figure 14: Empirical Hutter++ Model. Same setting as in Figure 6. Initial model trained on $T_{0}=100,000$ samples. No top-p inference or temperature scaling is used. $\beta=3/2$. In this setting, there is mild model collapse coming from the finite sample bias as well.
Figure 15: Empirical Hutter++ Model. Same setting as in Figure 14 with top $p=0.90$ synthesizing. No temperature scaling is used. $\beta=3/2$. Top-p selection significantly deteriorate the model collapse.
Figure 15: Empirical Hutter++ Model. Same setting as in Figure 14 with top $p=0.90$ synthesizing. No temperature scaling is used. $\beta=3/2$. Top-p selection significantly deteriorate the model collapse.
Figure 16: Empirical Hutter++ Model. Same setting as in Figure 14 with temperature 0.9 synthesizing. No top-p selection is used. $\beta=3/2$. Compared with Figure 14, temperature also create strong model collapse across multiple generation.
We now study mixing of clean and synthesized data in the bigram setting. Figures 18 and 18 add top-p tail-cutting when synthesizing, and start with $T_{0}=10,000$ original data samples, which are successively blended with synthesized data from the largest model. Note that in this setting we observe a reversion of scaling laws with increased AI data. This needs to be compared with the orange curve in Figure 20 in the deterministic Hutter setting. The probabilistic nature of the bigram models leads to a new effect here.
Figure 17: Empirical Hutter++ Model with Mixing. The initial ”clean” dataset comprises $T_{0}=10,000$ samples. For future generations, the largest model is used to synthesize data. For $T\leq 20,000$, training data is an equal mix of clean and generated data, for $T>20,000$ all clean data is used; the remaining training data is synthetic (so the ratio of clean data diminishes). Top-p 0.9, no temperature scaling, $\beta=3/2$.
Figure 17: Empirical Hutter++ Model with Mixing. The initial ”clean” dataset comprises $T_{0}=10,000$ samples. For future generations, the largest model is used to synthesize data. For $T\leq 20,000$, training data is an equal mix of clean and generated data, for $T>20,000$ all clean data is used; the remaining training data is synthetic (so the ratio of clean data diminishes). Top-p 0.9, no temperature scaling, $\beta=3/2$.
Figure 18: Empirical Hutter++ Model with Mixing. Same setting as in Figure 18 with top p 90, no temperature scaling, and $\beta=3/2$.
@@ Appendix C Proofs for the infinite memory (Hutter) Model (Sections 2 and 3)
Observe that the model $\widehat{f}$ makes an error on $i$ if and only if the $i$th ”skill” never occurred in the training dataset $\mathcal{D}_{T}$, i.e either (1) $i\geq k+1$, or (2) $1\leq i\leq k$ and $i_{t}\neq i$ for all $t\in[T]$. We deduce that
@@ Appendix F Details and Results from the Autoregressive Bigram model with Perplexity
We showcase experiments in the autoregressive bigram model with perplexity loss. We generate sequences of length $100$. Figures 23, Figure 23 and 25 aim to reproduce the ”paired bigram” Figure 12 in this setting, adding a top p mechanism and a temperature mechanism. Figure 27, Figure 27 and Figure 27 regenerates the setting of Figure 6 with the same top p and temperature.
Figure 22: Autoregressive Bigram Model with Perplexity Loss. Empirical confirmation of Theorem 2.4 for autoregressive data with top-p 1, Temperature 1. Each sequence data have length 100. Same setting as Figure 12. $\beta=3/2$.
Figure 22: Autoregressive Bigram Model with Perplexity Loss. Empirical confirmation of Theorem 2.4 for autoregressive data with top-p 1, Temperature 1. Each sequence data have length 100. Same setting as Figure 12. $\beta=3/2$.
Figure 23: Autoregressive Bigram Model with Perplexity Loss. Empirical confirmation of Theorem 2.4 for autoregressive data with top-p 0.9, Temperature 1. Each sequence data have length 100. $\beta=3/2$.
Figure 24: Autoregressive Bigram Model with Perplexity Loss. Empirical confirmation of Theorem 2.4 for autoregressive data with top-p 1, Temperature 0.9. Each sequence data have length 100. Same setting as Figure 12. $\beta=3/2$.
Figure 24: Autoregressive Bigram Model with Perplexity Loss. Empirical confirmation of Theorem 2.4 for autoregressive data with top-p 1, Temperature 0.9. Each sequence data have length 100. Same setting as Figure 12. $\beta=3/2$.
Figure 25: Autoregressive Bigram Model with Perplexity Loss. Each sequence data have length 100. Initial model trained on $T_{0}=10,000$ samples. It generates $T$ samples for Gen 1. Starting from Gen 2 models are trained on data generated by the most powerful model from the previous generation. Top-p 1, temperature 1, $\beta=3/2$.
Figure 26: Autoregressive Bigram Model with Perplexity Loss. Each sequence data have length 100. Same setting as Figure 25. Top-p 0.9, temperature 1, $\beta=3/2$.
Figure 26: Autoregressive Bigram Model with Perplexity Loss. Each sequence data have length 100. Same setting as Figure 25. Top-p 0.9, temperature 1, $\beta=3/2$.
Figure 27: Autoregressive Bigram Model with Perplexity Loss. Each sequence data have length 100. Same setting as Figure 25. Top-p 1, temperature 0.9, $\beta=3/2$.
@@ Appendix H Details of Experiments with Llama2
Utilizing one of the most advanced open-source models, Llama-2 7B, our research investigates the effects on LLMs when they undergo finetuning [^8] Quoting (Shumailov et al., 2023), we state that one can, in principle, replicate an experiment described here with training an LLM from scratch to demonstrate scaling law decay. Given that training a single moderately large model produces twice the American lifetime worth of CO2 (Strubell et al., 2019), we opted to not run such an experiment and instead focus on a more feasible finetuning setting. Note that just the language experiments described in the paper took weeks to run. with data generated by other LLMs. To ensure the generation of high-quality data and to provide a relevant but not trivial downstream task, we employ the Wikitext-103 dataset. We segment this dataset into chunks of 128 tokens, between each with a stride of 64 tokens, resulting in approximately 2.2 million chunks. Denote this dataset as $\mathcal{D}_{0}$. The task for generation involves producing the final 32 tokens given the initial 96 tokens from each chunk in the original dataset. In the initial generation (0-th generation), we use the Llama-2 7B FT model, which has been finetuned on $\mathcal{D}_{0}$, applying a generation loss that focuses solely on the cross-entropy loss of the final 32 tokens. We denote this initial model as $\mathcal{M}_{0}$, which demonstrates enhanced capacity for the generation task compared to the standard Llama-2 7B model. By querying $\mathcal{M}_{0}$ with the original 96 tokens from $\mathcal{D}_{0}$, we generate the dataset $\mathcal{D}_{1}$ and subsequently finetune Llama-2 7B on this dataset to obtain $\mathcal{M}_{1}$. This process is sequentially repeated to generate $\mathcal{D}_{i}$ from $\mathcal{M}_{i-1}$ and obtain $\mathcal{M}_{i}$ through finetuning. By comparing the performance of various $\mathcal{M}$ models on the test set derived from Wikitext-103, also segmented into 128-token chunks, we aim to investigate the model collapse in LLMs.
@@
Figure 34: Mixing Llama Generated Data with Original Data. Similar setting as Figure 4 left. Starting from gen 1, we mix the generated data with the original one with a ratio of 90 to 10. Top p 0.9 and temperature 0.9.
@@ Appendix I More Studies on Tail Cutting and Tail Narrowing Effects
Figure 35: Sequential bigram data: top p = 0.95 leads to similar effect as tail narrowing. 1000 data with sequence length 100.

### v2

@@ Appendix A Prior Work
demonstrating “emergence” of good BLEU performance with scale. This type of “emergence” has been massively confirmed in (Wei et al., 2022), where a working definition of “emerging” is “not present in smaller models, but appears in larger models”. In this sense, Wei et al. (2022) demonstrate empirically a large number of “skills” appearing with scale, like Multi-Task NLU, Modular arithmetic, word unscrambling and transliteration.
@@
Strengthening the tie between scaling laws and emergent skill, albeit in the opposite direction, Michaud et al. (2023) posit that skills that emerge in “quanta” imply a scaling law of the loss. Related, Chen et al. (2023) assume a hierarchy of skills to derive data curation mechanisms to precipitate the emergence of skills, though they do not allude to scaling laws directly.
@@ Appendix B Complimentary Figures for Sections 2, 3 and 4
Both Figures 3 (middle) and 12 illustrate the setting where each model consumes as much training data as its predecessor ($T_{0}=T$). We now relax the assumption that each successive model has strictly the same amount of training data as its predecessor. We assume that the generation 0 model is trained on $T_{0}$ (here, $T_{0}=100,000$) amount of original data to generate AI data for generation 1. All future generations, starting from generation 2, are trained on data generated by the most powerful model from the previous generation ($T=1,000,000$ data in this case). Figure 14 (for Hutter LLM) and 14 (for Hutter++ on paired bigram data) show the resulting scaling behavior. We take this setting even further by adding a top-$p^{inf}$ tail cutting mechanism and a temperature scaling mechanism for each synthetic data generation. Figure 6 cuts at $p=0.95$ and Figure 16 at temperature $0.9$.
@@
Figure 14: Empirical Hutter++ Model. Same setting as in Figure 6. Initial model trained on $T_{0}=100,000$ samples. No top-$p^{inf}$ inference or temperature scaling is used. $\beta=3/2$. In this setting, there is mild model collapse coming from the finite sample bias as well.
Figure 15: Empirical Hutter++ Model. Same setting as in Figure 14 with top $p^{inf}=0.9$ synthesizing. No temperature scaling is used. $\beta=3/2$. Top-$p^{inf}$ selection significantly deteriorate the model collapse.
Figure 16: Empirical Hutter++ Model. Same setting as in Figure 14 with temperature $\tau=$0.9 synthesizing. No top-$p^{inf}$ selection is used. $\beta=3/2$. Compared with Figure 14, temperature also create strong model collapse across multiple generation.
We now study mixing of clean and synthesized data in the bigram setting. Figures 18 and 18 add top-$p^{inf}$ tail-cutting when synthesizing, and start with $T_{0}=10,000$ original data samples, which are successively blended with synthesized data from the largest model. Note that in this setting we observe a reversion of scaling laws with increased AI data. This needs to be compared with the orange curve in Figure 20 in the deterministic Hutter setting. The probabilistic nature of the bigram models leads to a new effect here.
Figure 17: Empirical Hutter++ Model with Mixing. The initial “clean” dataset comprises $T_{0}=10,000$ samples. For future generations, the largest model is used to synthesize data. For $T\leq 20,000$, training data is an equal mix of clean and generated data, for $T>20,000$ all clean data is used; the remaining training data is synthetic (so the ratio of clean data diminishes). Top-$p^{inf}=$0.9, no temperature scaling, $\beta=3/2$.
Figure 18: Empirical Hutter++ Model with Mixing. Same setting as in Figure 18 with top-$p^{inf}=0.9$, no temperature scaling, and $\beta=3/2$.
@@ Appendix C Proofs for the infinite memory (Hutter) Model (Sections 2 and 3)
Observe that the model $\widehat{f}$ makes an error on $i$ if and only if the $i$th “skill” never occurred in the training dataset $\mathcal{D}_{T}$, i.e either (1) $i\geq k+1$, or (2) $1\leq i\leq k$ and $i_{t}\neq i$ for all $t\in[T]$. We deduce that
@@ Appendix F Details and Results from the Autoregressive Bigram model with Perplexity
We showcase experiments in the autoregressive bigram model with perplexity loss. We generate sequences of length $100$. Figures 23, Figure 23 and 25 aim to reproduce the ”paired bigram” Figure 12 in this setting, adding a top $p^{inf}$ mechanism and a temperature mechanism. Figure 27, Figure 27 and Figure 27 regenerates the setting of Figure 6 with the same top $p^{inf}$ and temperature.
Figure 22: Autoregressive Bigram Model with Perplexity Loss. Empirical confirmation of Theorem 2.4 for autoregressive data with top-$p^{inf}=$1, Temperature $\tau$ 1. Each sequence data have length 100. Same setting as Figure 12. $\beta=3/2$.
Figure 23: Autoregressive Bigram Model with Perplexity Loss. Empirical confirmation of Theorem 2.4 for autoregressive data with top-$p^{inf}=$0.9, Temperature $\tau$ 1. Each sequence data have length 100. $\beta=3/2$.
Figure 24: Autoregressive Bigram Model with Perplexity Loss. Empirical confirmation of Theorem 2.4 for autoregressive data with top-$p^{inf}=$1, Temperature $\tau=$0.9. Each sequence data have length 100. Same setting as Figure 12. $\beta=3/2$.
Figure 25: Autoregressive Bigram Model with Perplexity Loss. Each sequence data have length 100. Initial model trained on $T_{0}=10,000$ samples. It generates $T$ samples for Gen 1. Starting from Gen 2 models are trained on data generated by the most powerful model from the previous generation. Top-$p^{inf}=$1, temperature $\tau=$1, $\beta=3/2$.
Figure 26: Autoregressive Bigram Model with Perplexity Loss. Each sequence data have length 100. Same setting as Figure 25. Top-$p^{inf}=$0.9, temperature $\tau=$1, $\beta=3/2$.
Figure 27: Autoregressive Bigram Model with Perplexity Loss. Each sequence data have length 100. Same setting as Figure 25. Top-$p^{inf}=$1, temperature $\tau=$0.9, $\beta=3/2$.
@@ Appendix H Details of Experiments with Llama2
Utilizing one of the most advanced open-source models, Llama-2 7B, our research investigates the effects on LLMs when they undergo finetuning [^8] One can, in principle, replicate an experiment described here with training an LLM from scratch to demonstrate scaling law decay. We opted to not run such an experiment and instead focus on a more feasible finetuning setting, since just the language experiments described in the paper took weeks to run. with data generated by other LLMs. To ensure the generation of high-quality data and to provide a relevant but not trivial downstream task, we employ the Wikitext-103 dataset. We segment this dataset into chunks of 128 tokens, between each with a stride of 64 tokens, resulting in approximately 2.2 million chunks. Denote this dataset as $\mathcal{D}_{0}$. The task for generation involves producing the final 32 tokens given the initial 96 tokens from each chunk in the original dataset. In the initial generation (0-th generation), we use the Llama-2 7B FT model, which has been finetuned on $\mathcal{D}_{0}$, applying a generation loss that focuses solely on the cross-entropy loss of the final 32 tokens. We denote this initial model as $\mathcal{M}_{0}$, which demonstrates enhanced capacity for the generation task compared to the standard Llama-2 7B model. By querying $\mathcal{M}_{0}$ with the original 96 tokens from $\mathcal{D}_{0}$, we generate the dataset $\mathcal{D}_{1}$ and subsequently finetune Llama-2 7B on this dataset to obtain $\mathcal{M}_{1}$. This process is sequentially repeated to generate $\mathcal{D}_{i}$ from $\mathcal{M}_{i-1}$ and obtain $\mathcal{M}_{i}$ through finetuning. By comparing the performance of various $\mathcal{M}$ models on the test set derived from Wikitext-103, also segmented into 128-token chunks, we aim to investigate the model collapse in LLMs.
@@
Figure 34: Mixing Llama Generated Data with Original Data. Similar setting as Figure 4 left. Starting from gen 1, we mix the generated data with the original one with a ratio of 90 to 10. Top-$p^{inf}=$0.9 and temperature $\tau=$0.9.
@@ Appendix I More Studies on Tail Cutting and Tail Narrowing Effects
Figure 35: Sequential bigram data: top $p^{inf}$ = 0.95 leads to similar effect as tail narrowing. 1000 data with sequence length 100.